OpenStudy (anonymous):

The first term of a geometric sequence is 64, and the second term is 16. Find the fifth term. Can someone help me?

OpenStudy (anonymous):

Call the terms: \[a + ar + ar^2 + .... + ar^{n-1} \] a = 64 ar = 16 => r = ... n = 5 So last term is \[ ar^{5-1} = ar^5 = ...\]

OpenStudy (anonymous):

so the answer is 4000?